OpenStudy (anonymous):

Maurice directs a beam of light on a sheet of glass at an angle of 51°. What is the angle of refraction in the glass? The refractive index of glass is 1.46.

OpenStudy (anonymous):

Snell's law is \[n_1 \sin\theta_1 = n_2 \sin\theta_2\] Assume the refractive index of air is 1 \[\sin 51^\circ = 1.46\sin\theta_2\]